OpenStudy (anonymous):

What are the x-coordinates for the maximum points in the function f(x) = 4 cos(2x − π) from x = 0 to x = 2π?

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

sorry mate no idea..

OpenStudy (anonymous):

@aum @Loser66 @LostInTheMusic

OpenStudy (anonymous):

I got the answer I just want to know if it is correct

OpenStudy (aum):

In the interval \([0, 2\pi]\), \(\cos(\theta)\) is maximum when \(\theta = 0, 2\pi\). Therefore, \(\cos(2x - \pi)\) is maximum when \(2x - \pi = 0, 2\pi\) or \(x = \pi/2,~~x= 3\pi/2\)

OpenStudy (anonymous):

Yes I got it correct